OpenStudy (anonymous):

What are the next three terms of the sequence 1, -3, 9, -27

OpenStudy (ddcamp):

Each term is the previous term × -3 -27 × -3 = 81 81 × -3 = -243 -243 × 3 = 729

OpenStudy (anonymous):

by looking at the equation, the pattern is \[(-3)^{n}\] where n=(0,1,2,3,......) so, the next 3 terms will be \[(-3)^{4}=81\]\[(-3)^{5}=-243\]\[(-3)^{6}=729\]

OpenStudy (chrisplusian):

I would have come up with the same as @shrinifores